# A Black Hole Airy Tail

### v3

@@ V Discussion
We showed how wormholes in JT gravity restore positivity of $S_{SQ}^{(n)}$. This required a resummation of the genus expansion, and no single geometry was sufficient. This suggests that requiring positivity of $S_{SQ}^{(n)}$ can put constraints on wormhole contributions in a putative gravitational theory: a wormhole bootstrap.
We probed a regime where all the entropies are non-self-averaging and thus are not necessarily representative of a typical member of the ensemble. Both $S_{SQ}^{(n)}$ and $S_{Q}^{(n)}$ in this regime are dominated by atypical members of the ensemble. This is most obvious for $S_{SQ}^{(n)}$ from the one-eigenvalue instanton analysis, which is a highly atypical member of the ensemble with a large gap between the two smallest eigenvalues. Thus, our results do not carry forward in an obvious manner to the gravitational path integral in higher dimensions if it is not dual to an ensemble of theories.
Finally, we note that $S_{A}^{(n)}$ is the simplest quantity, requiring only the average density of states. Our newly defined $S_{SQ}^{(n)}$ is close behind since it only requires up to the $n$th moment of the density of states. On the other hand, $S_{Q}^{(n)}$ requires arbitrarily many moments of the density of states. This is the key reason for the subtlety in performing the replica limit required.
@@ Acknowledgements.
Appendix A Semi-quenched Rényi entropy at higher $n$
@@
with $\alpha_{m+1}\equiv\alpha_{1}$. Next, by symmetrizing this function over the arguments:
$\mathcal{E}^{(m)}_{\text{sym}}(\alpha_{1},\dots,\alpha_{m})\equiv\frac{1}{m}\sum_{\sigma}\mathcal{E}^{(m)}(\alpha_{\sigma(1)},\dots,\alpha_{\sigma(m)}).$ (25)
@@
$\langle Z(\beta)^{n}\rangle=\sum_{\pi\in\Pi_{n}}(-1)^{\ell(\pi)+1}\mathcal{E}_{\text{sym}}^{(\ell(\pi))}(\alpha_{\pi})\evaluated{}_{\alpha_{i}=\alpha}.$ (26)
@@
Appendix B Instanton in JT Gravity
@@
Appendix C Two-eigenvalue instanton
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v4

@@ V Discussion
We showed how wormholes in JT gravity restore positivity of $S_{SQ}^{(n)}$, which is sufficient to prove that the ground state is isolated and the quenched entropy also remains positive. This required a resummation of the genus expansion, and no single geometry was sufficient. Computing other observables sensitive to the isolated ground state is now also possible: they will also include the same genus resummation. Our results suggest that requiring positivity of entropies can put constraints on wormhole contributions in a putative gravitational theory: a wormhole bootstrap.
We probed a regime where all the entropies are non-self-averaging and thus are not necessarily representative of a typical member of the ensemble. Both $S_{SQ}^{(n)}$ and $S_{Q}^{(n)}$ in this regime are dominated by atypical members of the ensemble. This is most obvious for $S_{SQ}^{(n)}$ from the one-eigenvalue instanton analysis, which is a highly atypical member of the ensemble with a large gap between the two smallest eigenvalues. Thus, our results for the entropies are not as valuable in higher dimensions if gravity is not dual to an ensemble of theories. Nonetheless, the positivity of the semi-quenched entropy is a consequence of the presence of an isolated ground state in every member of the ensemble. This is a self-averaging property that is meaningful in higher dimensions.
@@ Acknowledgements.
Appendix A Isolated ground state from semi-quenched positivity
The goal of this appendix is to prove that if even a single member of an ensemble has a continuous density of states (DOS) including the ground state, then the semi-quenched entropy becomes negative at sufficiently small temperature. As mentioned in the main text, this only happens if,
$\langle Z_{n}\rangle>\langle Z_{1}^{n}\rangle.$ (24)
Let us assume there exists at least one member of the ensemble that has a continuous DOS above the ground state with a Laurent expansion
$\rho(E)=e^{S_{\text{low T}}}(E-E_{0})^{\alpha}+o((E-E_{0})^{\alpha})$ (25)
where $e^{S_{\text{low T}}}$ keeps track of the overall scaling of the density of states at low temperatures and $\alpha>-1$ (due to the convergence of the partition function) determines the behavior of the DOS near the edge. For such an ensemble member (24) becomes
$\beta>e^{\frac{S_{\text{low T}}}{1+\alpha}}n^{\frac{1}{n-1}}\Gamma(1+\alpha)^{\frac{1}{1+\alpha}}\,.$ (26)
Thus, there always exists some sufficiently large value of $\beta$ for which the inequality (24) is satisfied for this ensemble member. For ensemble members that have an isolated ground state, the expansion (25) is invalid and $Z_{1}^{n}\to Z_{n}$ as $\beta\to\infty$, with $Z_{1}^{n}\leq Z_{n}$ for all $\beta$. Thus, averaging over an ensemble where some members have an isolated ground state and some do not, the inequality (24) is satisfied, and the semi-quenched entropy becomes negative. The positivity of the semi-quenched entropy even for fixed $n$ but all $\beta$ thus implies that all members of the ensemble have an isolated ground state.
Appendix B Semi-quenched Rényi entropy at higher $n$
@@
with $\alpha_{m+1}\equiv\alpha_{1}$. Next, by symmetrizing this function over the arguments,
$\mathcal{E}^{(m)}_{\text{sym}}(\alpha_{1},\dots,\alpha_{m})\equiv\frac{1}{m}\sum_{\sigma}\mathcal{E}^{(m)}(\alpha_{\sigma(1)},\dots,\alpha_{\sigma(m)}),$ (28)
@@
[…]
@@
Appendix C Instanton in JT Gravity
@@
Appendix D Two-eigenvalue instanton
@@
$=N\int d\lambda\,p\left(\lambda\right)e^{-\beta\lambda},$ (35)
@@
$\langle Z\left(\beta\right)^{2}\rangle\supset N\int d\lambda\,p\left(\lambda\right)e^{-2\beta\lambda},$ (39)
@@
$\langle Z(\beta)^{2}\rangle\supset N(N-1)\int d\lambda_{1}d\lambda_{2}\,p\left(\lambda_{1},\lambda_{2}\right)e^{-\beta\left(\lambda_{1}+\lambda_{2}\right)},$ (40)
@@
$\int d\lambda_{1}d\lambda_{2}\,p\left(\lambda_{1},\lambda_{2}\right)e^{-\beta\left(\lambda_{1}+\lambda_{2}\right)}\approx e^{\frac{\alpha^{3}}{12\gamma^{3}}},$ (41)